### v1

# ProtChatGPT: Towards Understanding Proteins with Large Language Models

Abstract
Protein research is crucial in various fundamental disciplines, but understanding their intricate structure-function relationships remains challenging. Recent Large Language Models (LLMs) have made significant strides in comprehending task-specific knowledge, suggesting the potential for ChatGPT-like systems specialized in protein to facilitate basic research. In this work, we introduce ProtChatGPT, which aims at learning and understanding protein structures via natural languages. ProtChatGPT enables users to upload proteins, ask questions, and engage in interactive conversations to produce comprehensive answers. The system comprises protein encoders, a Protein-Language Pertaining Transformer (PLP-former), a projection adapter, and an LLM. The protein first undergoes protein encoders and PLP-former to produce protein embeddings, which are then projected by the adapter to conform with the LLM. The LLM finally combines user questions with projected embeddings to generate informative answers. Experiments show that ProtChatGPT can produce promising responses to proteins and their corresponding questions. We hope that ProtChatGPT could form the basis for further exploration and application in protein research. Code and our pre-trained model will be publicly available.
Protein Understanding, Large Language Models, ChatGPT-like system
1 Introduction
Proteins, as essential molecular entities for life, hold paramount significance in biological processes. The comprehensive understanding of protein structure and function is of utmost importance for advancing research in the realms of biology and biomedicine. However, traditional protein research normally involves labor-intensive laboratory experiments and extensive literature reviews, which could be time-consuming and require specialized expertise in protein.
Recently, Large Language Models (LLMs), e.g., ChatGPT (Radford et al., 2019), have prevailed in Natural Language Processing (NLP) (Devlin et al., 2018; Raffel et al., 2020; Touvron et al., 2023; Chowdhery et al., 2022). With superior language understanding and logical reasoning capabilities, these models can perform various intricate linguistic tasks such as question and answering (Q&A). Considering that protein sequences can be regarded as their own “natural language”, this intuitively motivates us to ride on LLMs’ coattails and customize them into protein research based on large-scale biological corpora, e.g., RCSB-PDB (Guo et al., 2023).
Empirically, with the capabilities of LLMs specialized in protein, researchers can potentially achieve (1) Protein Understanding and Analysis by simplifying the retrieval of crucial information (e.g., structures, functions, interactions, mutations, and disease associations) about specific proteins for research; (2) Customized Protein Design by characterizing the patient’s unique protein structures to discover targeted drugs and further verify expected functions for healthcare.
In this paper, we propose an AI-based protein chat system, named ProtChatGPT, to implement ChatGPT-like functionalities for the protein research field. ProtChatGPT works in a similar principle with natural language conversation systems. Users are allowed to upload protein 1D sequences or 3D structures (e.g., fasta or pdb files) and pose diverse related questions. Then, ProtChatGPT produces comprehensive responses in an interactive manner based on the questions. In this way, researchers can intuitively acquire valuable insights and interactively delve into the complexities of diverse proteins. Specifically, ProtChatGPT consists of four components: protein encoders, a Protein-Language Pre-training Transformer (PLP-former), a projection adapter and an LLM, as shown in Figure 1. First, we employ two pre-trained protein encoders to embed the 1D (i.e., by ESM-1b (Rives et al., 2021) ) and 3D structures (i.e., by ESM-IF1 (Hsu et al., 2022)), respectively. Then, to align the protein and language modalities, we propose the PLP-former. PLP-former extracts features from the output of the protein encoder, and learns the protein representations that are most relevant to the text description. Third, we use an adapter as an information transmitter to convert protein embeddings into protein prompts that can be interpreted by the LLM. Finally, the LLM combines user questions (i.e., question prompts) with the transmitted protein prompts to produce corresponding answers. We conduct experiments on protein understanding and design. Experimental results show the effectiveness of the proposed method. In summary, our contributions are as follows:
  • We propose ProtChatGPT, an interactive ChatGPT-like system that engages Q&A for protein-related research, which significantly facilitates protein understanding and design.
  • We introduce PLP-former, a transformer-based module that aligns the protein with its corresponding description.
  • We propose a two-stage strategy that bootstraps protein-language pre-training from off-the-shelf pre-trained protein encoders and frozen large language models.
  • We demonstrate ProtChatGPT’s versatility and range of applications by deploying it to tasks of a rather distinct nature, including protein understanding and design.
Figure 1: An overview of the ProtChatGPT framework. The training process consists of two stages: (1) protein-description representation learning stage, and (2) protein-to-text generative learning stage. In the first stage, we enforce the PLP Transformer, a lightweight transformer with learnable query tokens, to extract features from a frozen protein 1D encoder, and learn the protein representation most relevant to the text description. The PLP-former takes sequence-description pairs as input, and outputs the learned tokens as the sequence embedding. In the second stage, we perform protein-to-text generative learning by connecting the output of the PLP Transformer as well as a frozen 3D encoder to a frozen LLM, and trains a multi-level adapter as an information bottleneck between two stages, such that its output protein representation can be interpreted by the LLMs. Finally, the LLM can produce descriptive answers given the question prompt and the multi-level (sequence and structure) protein prompt from the adapter.
2 Related Work
Protein Representation Learning. Proteins are workhorses of the cell, which contain four distinct levels of structures carrying out their fundamental functions. Previous protein representation works seek to learn protein representations based on different levels of proteins. Considering protein sequences as language in life, several works (Madani et al., 2023; Notin et al., 2022) encode amino acid tokens using the Transformer model (Vaswani et al., 2017) to extract pairwise relationships among amino acids, and autoregressively recover protein sequences on extensive protein sequence databases. Alternatively, other sequence modeling methods(Lin et al., 2023; Meier et al., 2021; Rives et al., 2021; Rao et al., 2020; Vig et al., 2020) resort to use Masked Language Modeling (MLM) to develop attention patterns that correspond to the residue-residue contact map of the protein. Compared with sequence-based methods, structure-based methods (Gligorijević et al., 2021; OpenAI, 2023; Zhang et al., 2023) directly dictate protein functions and encode geometric information of proteins for topology-sensitive tasks such as molecule binding (Jin et al., 2021; Kong et al., 2022), protein interface analysis (Mahbub & Bayzid, 2022; Réau et al., 2023), and protein properties prediction (Zhang et al., 2022). In this paper, we aim to leverage these pre-trained Large Protein Models (LPMs) for high-quality embeddings without fine-tuning their network parameters.
Large Language Models (LLMs). Recently, Natural Language Processing (NLP) has witnessed significant advances due to the development of Large Language Models (LLMs) (Radford et al., 2019; Devlin et al., 2018) trained on an extensive, diverse corpus. Consequently, many multi-modal variants (Taylor et al., 2022; Alayrac et al., 2022; Jing et al., 2020; Wang et al., 2023; van Sonsbeek et al., 2023) based on LLMs have gained significant attention for the understanding of information in other modalities beyond text. For example, BLIP-2(Li et al., 2023) designs a Q-Former to align the visual features from the frozen visual encoder with large language models. FROMAGe (Koh et al., 2023) freezes the LLM and visual encoders, and fine-tunes linear mapping layers to achieve cross-modality interactions. Similarly, MedVQA (van Sonsbeek et al., 2023) employs a multi-layer perceptron (MLP) network that maps the extracted visual features from a frozen vision encoder to a set of learnable tokens, which develops an open-ended VQA for diagnoses and treatment decisions. Galactica (Taylor et al., 2022) explicitly models the protein sequences and SMILES with scientific literature, and enables the model to explain the properties of the sequences. ProteinChat (Guo et al., 2023) further models the protein structure with its corresponding descriptions in Protein Data Bank for the protein Q&A task. However, this method only considers the impact of protein structure on its function, while neglecting the complementarity between different levels of structure, such as amino acid sequences (primary structure) and geometric coordinates (tertiary structure). In this paper, we aim to adapt pre-trained general LLMs (e.g., Vicuna (Chiang et al., 2023)) for protein-specific ChatGPT-like tasks, which aligns multi-level protein features from LPMs with LLMs.
Vision-Language Pretraining (VLP). Data collected from different modalities generally offer distinct perspectives, frequently synergizing to yield a comprehensive understanding, enhancing the overall comprehension of the data. Vision-language pre-training (VLP) aims to learn multimodal foundation models, showing improved performance on various vision-and-language tasks (Radford et al., 2021a), Existing VLP methods can be roughly divided into representation learning-based and generative learning-based. Representation learning-based methods (Radford et al., 2021b; Jia et al., 2021; Yao et al., 2021; Li et al., 2022b, 2021) usually consider the image-text pairs as multi-modal views of the same semantics, and perform contrastive or multi-view learning for the alignment between multiple modalities. Generative learning-based methods  (Li et al., 2019; Lu et al., 2019; Chen et al., 2020; Li et al., 2020; Zhang et al., 2021; Wang et al., 2022; Zeng et al., 2021; Bao et al., 2022) aim to reconstruct the corrupted text (image) with the assistance of visual (text) modality through MLM-like objectives. For example, SimVLM (Wang et al., 2021) introduces a single prefix language modeling (PrefixLM) objective for exploiting large-scale weak supervision in VLP. CoCa (Yu et al., 2022) further verifies the representation ability of autoregressive language modeling in the vision-language domain. In this paper, we consider protein as a specialized biological language that encodes and communicates biological information through its amino acid sequences and interactions. Inspired by existing VLP methods (Li et al., 2023; Zhu et al., 2023), we first propose a representation learning-based Protein-Language Pre-training (PLP) framework to understand protein sequences via natural language instructions, and then design a generative learning-based multi-level adapter to generate the question-related answers combined with extra structural embeddings.
3 Methods
While demonstrating excellent performance in natural language tasks, LLM cannot directly facilitate protein question-answering tasks due to the modality gap between protein structures and biomedical texts. In order to bridge this gap, we introduce a protein-language pre-training strategy with two stages: (1) protein-description representation learning stage and (2) protein-to-text generative learning stage.
3.1 Architecture
As shown in Figure 1, ProtChatGPT consists of two pre-trained protein encoders, a trainable PLP-former, a trainable multi-level projection adapter and a pre-trained LLM decoder, working synergistically to provide protein-related insights.
Multi-Level Protein Encoders
First, we use a pre-trained sequence encoder to extract protein sequence features. Specifically, given a protein sequence with $N$ amino acids, the encoder produces the corresponding sequence embedding $\boldsymbol{E}_{seq}\in\mathbb{R}^{N\times C_{seq}}$, where $C_{seq}$ is the number of embedding channels. In our implementation, we use ESM-1b (Rives et al., 2021) as the 1D sequence encoder, where $C_{seq}$ = 768. Although ESM-1b is able to implicitly capture structural contact information, incorporating detailed 3D structures explicitly can be an effective way to model spatial interactions between residues. Therefore, we further propose to enhance the ESM-1b with a supplementary protein structure encoder ESM-IF1 (Hsu et al., 2022). Specifically, we select the feature from an intermediate layer as a complementary structure embedding $\boldsymbol{E}_{str}\in\mathbb{R}^{N\times C_{str}}$ where the number of embedding channels $C_{str}=512$, indicating the geometric protein knowledge. By acquiring such multi-level property information, the quality of protein embeddings can be further improved, considering that the protein properties studied in ESM-1b and ESM-IF1 can correlate with each other. Note that, both the sequence and structure encoder are frozen for efficient training.
Figure 2: Illustrations of the PLP-Former and protein-language representation learning. PLP-Former consists of two transformer submodules with shared self-attention: (1) a text transformer that performs encoding and decoding of protein descriptions, and (2) a protein transformer that interacts with the frozen ESM-1b for sequence feature extraction. PLP-former is trained by jointly optimizing three pre-training objectives (dashed boxes) on sequence-description pairs.
Protein-Language Pretraining Transformer (PLP-former).
As mentioned before, the key challenge of transferring LLMs to protein research lies in the modality gap between protein structures and biomedical texts. Despite the strong language generation and zero-shot transfer abilities of LLMs, directly retraining them end-to-end for protein specialization appears to be impractical due to the massive number of parameters and data requirements. Another alternative is fine-tuning the pre-trained parameters, but this often leads to catastrophic forgetting. Considering this trade-off, we propose a Protein-Language Pre-training Transformer (PLP-former) for efficient cross-modal alignment between protein and text, while remaining LLMs frozen during the training. Following existing vision-language works (Li et al., 2023; Zhu et al., 2023; Dai et al., 2023), we use the PLP-former to extract protein-related features from a frozen protein sequence encoder.
As shown in Figure 2, the input of PLP-former consists of three parts: sequence embedding from ESM-1b, the corresponding description, and a set of learnable tokens. The learnable tokens $\boldsymbol{T}$ first perform mutual interactions via self-attention layers in the protein transformer. Specifically, the tokens $\boldsymbol{T}$ first acquire queries $\mathbf{Q}$, keys $\mathbf{K}$ and values $\mathbf{V}$ through three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$ and $\mathbf{W}_{v}$. It can be formulated as:
$\mathbf{Q}=\boldsymbol{E}_{str}^{l-1}\mathbf{W}_{q},\mathbf{K}=\boldsymbol{E}_{seq}^{l-1}\mathbf{W}_{k},\mathbf{V}=\boldsymbol{E}_{seq}^{l-1}\mathbf{W}_{v}.$ (1)
Next, the attention map $w$ is computed by taking the dot product of Q and K as:
[…]
where $d_{k}$ represents the dimensionality of the keys, softmax is the softmax activation function. The refined tokens $\boldsymbol{T}^{\prime}$ of the self-attention block can be written as:
$\boldsymbol{T}^{\prime}=\mathbf{W}_{o}\times(w\cdot\mathbf{V}),$ (3)
where $\cdot$ represents the dot product, $\times$ means the matrix multiplication, and $\mathbf{W}_{o}$ is the projection matrix for output. Given the sequence features from ESM-1b as $\boldsymbol{E}_{seq}$, tokens then interact with $\boldsymbol{E}_{seq}$ through cross-attention layers, which can be formulated as:
[…]
where $\boldsymbol{T}^{\prime\prime}$ represents the refined tokens after cross-attention, $\mathbf{W}_{q}^{\prime}$, $\mathbf{W}_{k}^{\prime}$, $\mathbf{W}_{v}^{\prime}$ and $\mathbf{W}_{o}^{\prime}$ are a new set of learning transformation matrices. Additionally, benefiting from the text transformer, tokens can further interact with the textual descriptions through the same self-attention layers. Finally, the PLP-former produces the output $\boldsymbol{E}_{seq}$ after a linear feed-forward layer (Vaswani et al., 2017). Depending on the pre-training task, we implement distinct attention masking strategies (Li et al., 2023) within the self-attention block to regulate the token-text interaction. The training details of PLP-former are given in Appendix A. In this way, PLP-Former can effectively select the most useful information for the LLM while removing irrelevant protein information. This reduces the burden of the LLM to learn protein-language alignment, thus mitigating the catastrophic forgetting problem.
Note that we apply PLP only to the sequence embeddings since the reported protein structures are much less than sequences. For example, there are 182K experimentally-determined structures in Protein Data Bank (PDB) (Berman et al., 2000) while 47M protein sequences in Pfam (Mistry et al., 2021). Thus we only use selected sequence-structure pairs during the second-stage training.
Multi-Level Projection Adapter.
For the second protein-to-text generative learning stage, we further design a multi-level projection adapter to harvest the LLM’s generative language capability. The adapter takes the pre-aligned sequence embedding $\boldsymbol{E}_{seq}$ from PLP-former and structure embedding $\boldsymbol{E}_{str}$ from ESM-IF1 as inputs, and acts as an information bottleneck to the LLM decoder, such that its output protein representation can be interpreted by the LLM. In practice, we use two individual Fully-Connected (FC) layers to linearly project the output protein embeddings into the same dimension as the question embedding of the LLM. They function as soft protein prompts that condition the LLM on protein representation from 1D and 3D levels.
Large Language Models (LLMs).
Finally, the projected protein prompts are prepended to the question prompts (text embeddings of user questions) through concatenation. In implementation, we deploy the Vicuna-13b (Chiang et al., 2023) as our LLM decoder, which employs the Transformer decoder (Vaswani et al., 2017) to model the conditional generation probability $p_{\theta}(n_{i}|n_{<i})$ in the language model. Specifically, given the context vector of tokens $N_{i-1}$ (i.e., protein embeddings along with user questions), the generated tokens $n_{i}$ (i.e., answers) are computed as follows:
$h_{0}=N_{i-1}\mathbf{W}_{e}+\mathbf{W}_{p},N_{i-1}=(n_{1},n_{2},...,n_{i-1})$ (5)
$h_{l}=transformer(h_{l-1})(1\leq l\leq m)$
$p(n_{i}|N_{i-1})=softmax(h_{m}\mathbf{W}_{e}^{T})$
where $m$ refers to the number of layers, $W_{e}$ denotes the token embedding matrix, $W_{p}$ represents the positional embedding matrix, and $transformer$ is a decoder layer consisting of a multi-head self-attention block and a position-wise feedforward network. During the second training stage, the generated tokens $n_{i}$ are replaced with publicly available descriptions for protein-to-text generative learning, while remaining empty in the testing stage, expecting ProtChatGPT to generate descriptive answers for the given protein and corresponding questions.
3.2 Protein-Language Pre-training Datasets
In order to train ProtChatGPT, dedicated protein-specific training datasets are indispensable for our proposed two-stage training strategy. Details and examples of our training datasets are given in Appendix B. During the first stage, we adopt ProtDescribe dataset (Xu et al., 2023) to train the PLP-Transformer for protein-description representation learning. ProtDescribe dataset contains 553,052 aligned pairs of protein sequences and textual property descriptions such as protein names, functions, families, subcellular locations, etc. The whole dataset is used for the first-stage pre-training. After the first stage, PLP-Transformer is able to acquire the protein 1D representation that is most pertinent to the textual description.
Nevertheless, despite covering the protein sequence positions, ProtDescribe de facto simply relies on textual descriptions to provide a rough indication of protein structural and functional similarity. It might be more straightforward to directly utilize structural information, especially considering that ESM-IF is specifically designed for protein geometric structures. Considering this problem, we resort to the RCSB-PDB Protein Description Dataset (Guo et al., 2023) which comprises 143,508 structure-description aligned pairs of proteins. We further expand this dataset with 1D sequences, enabling the joint training with our PLP-former. For every protein taken into account, we compile its 1D residue sequences, 3D atomic coordinates, along with corresponding descriptions to train our multi-level adapter. We randomly select 1,000 pairs of protein for evaluation, and the rest part of the dataset is used for our second-stage pre-training.
3.3 Aligned Protein-Text Generation
To enhance the model training with protein-text pairs, we utilize a specialized token prompt, following the conversational format used in Vicuna-13b:
$\begin{array}[]{l}\mathbf{Q}{:}<Protein><ProteinPrompts></Protein>\\ <QuestionPrompts>\\ \mathbf{A}{:}<Descriptions>\end{array}$
[…]
4 Experiments
4.1 Implementation Details
For our training setup, we freeze both the sequence and structure protein encoders, as well as the LLM decoder, solely focusing on training the PLP-former (first stage) and the projection adapter (second stage). This two-stage strategy ensures that the pre-trained models retain their learned knowledge while fine-tuning the projection layer to align the protein embeddings with the LLM’s requirements. For the first stage, the PLP-former is initialized with the pre-trained weights of PubMedBERT (Gu et al., 2021), whereas the cross-attention layers and learnable tokens are randomly initialized. Following the same setting as PubMedBERT, we use 32 learnable tokens with a dimension of 768, which is the same as the hidden dimension of the PLP-Former. We pre-train the PLP-former on the ProtDescribe dataset for 20K epochs with a batch size of 64. We use the AdamW (Loshchilov & Hutter, 2017) optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.98, and a weight decay of 0.05. We use a cosine learning rate decay with a peak learning rate of 1e-4, a minimum learning rate of 8e-5, and a linear warm-up of 5K iterations. For the second stage, we freeze the PLP-former and LLM, and train the projection adapter for 1K epochs with a batch size of 128. The minimum learning rate for the second stage is set as 5e-5. To avoid out-of-memory issues, we randomly trim excessively long protein sequences and structures to 1000 amino acids. All experiments are performed on 4 NVIDIA A100 (80GB) GPUs, our model with ESM-1b and Vicuna-13b requires 5.5 days for the first stage and 2 days for the second stage.
4.2 Results
4.2.1 Qualitative Results
We first show some example conversations between the user and ProtChatGPT to indicate the promising results on both protein understanding and design tasks. For a fair comparison, we use the protein that does not appear in the training set of ProtChatGPT. Figure 3 shows some conversation examples of our ProtChatGPT related to protein understanding and design, as well as customized protein design tasks. It can be observed that our system comprehends the meaning of the question well, providing accurate and logically consistent responses that not only identify the types and characteristics of proteins, but also extend to exploring their related application areas and potential values. Additionally, when presented with a protein, ProtChatGPT can provide explanations regarding questions associated with protein understanding. This implies that researchers can swiftly survey the structures, functions, and other relevant properties through ProtChatGPT. Coupled with this comprehension capability, further judgments on mutations and disease association can be made about specific proteins, leading to the potential of targeted protein design in healthcare.
Figure 3: Dialogue examples of ProtChatGPT for protein understanding and analysis (a, b) as well as customized protein design (c) tasks.
Figure 4: Case studies on homologous proteins (a, b) and mutually exclusive functions (c).
4.3 Case Study
To further validate the utility of our method, we conduct case studies on two specific examples, including homologous proteins and mutually exclusive functions.
4.3.1 Homologous Proteins
Homologous proteins are proteins that are derived from the same ancestral gene, which usually share similar amino acid sequences and structures. This study could help in understanding evolutionary relationships, predicting protein functions, and identifying potential therapeutic interventions. In this case, we choose Actin and Myosin proteins for our test. As shown in Figure 4 $(a)$ and $(b)$, although these two proteins have similar amino acid sequences, our multi-level approach can still differentiate and make reasonable analyses due to their significant structural differences.
4.3.2 Mutually Exclusive Functions
Mutually exclusive protein functions describe situations where a certain protein in the same cell cannot simultaneously perform its roles, especially in signaling pathways, regulatory mechanisms, or cellular processes, where one activity inhibits or prevents another. Understanding these functions is important for grasping how cells make decisions and respond to their environment, particularly in complex processes like development, immune responses, and disease progression. In this case, we choose the p53 protein for our test. Figure 4 (c) indicates that our method can efficiently incorporate contextual semantics to provide varied analyses under different environment prompts.
4.3.3 Quantitative Comparisons
To validate the effectiveness of our method, we further conduct several quantitative experiments. To fully showcase the capability of ProtChatGPT, we randomly select 1,000 protein sequence-structure-description pairs from the RCSB-PDB Protein Description Dataset dataset to serve as the test set. Note that these testing protein pairs are not used during training for a fair comparison. We employ seven commonly used metrics in the image captioning and NLP domains to test the performance of ProtChatGPT. Detailed descriptions of these metrics can be found in Appendix C.
We first devise two variants to validate our contribution. (1) w/o structure: remove the supplementary branch of the 3D structure encoder, relying solely on sequence information for learning. (2) w/o PLP-former: remove the PLP-former and directly aligned the 1D and 3D embeddings using the adapter. As a fair comparison, we give the LLM with the same user question as: “Describe this protein in short”. It can be observed from Table 1 that both the structural encoder and the PLP-former play indispensable roles in supplementation and alignment, respectively. Notably, in contrast to common metrics like BLEU and METEOR, SPICE and BERTScore pay more attention to deeper semantic information, rather than just lexical and syntactic alignment. Particularly, we replace the original Bert encoder with PubMedBERT (Gu et al., 2021), a biomedical description-specific encoder pre-trained on large-scale datasets. To some extent, this indicator can reflect the scientific validity of the generated responses in the biomedicine domain.
Table 1: Quantitative Comparisons on the proposed multi-level encoders and PLP-former. $\uparrow$ indicates that a higher value corresponds to better performance. The best performances are marked in bold.
VariantMetric BLEU-1 normal-$\uparrow$ BLEU-4 normal-$\uparrow$ ROUGE-L normal-$\uparrow$ METEOR normal-$\uparrow$ CIDEr normal-$\uparrow$ SPICE normal-$\uparrow$
PubMed
BERTScore normal-$\uparrow$
w/o structure 0.457 0.311 0.405 0.237 0.504 0.231 0.335
w/o PLP-former 0.581 0.352 0.463 0.270 0.572 0.276 0.421
ProtChatGPT 0.610 0.394 0.489 0.291 0.638 0.316 0.457
Figure 5: Comparison of fine-tuning of PLP-former and LLM decoder during the second-stage training. We compute the SPICE and PubMed BERTScore for semantic evaluation.
Furthermore, we modified the two-stage training strategy of PLP in an attempt to achieve better protein-specific dialogue capabilities. As shown in Figure 5, we independently fine-tune the LLM decoder and PLP-former in the second stage with the RCSB-PDB dataset. The performance on two high-level semantic metrics SPICE and PubMed BERTScore indicate that further fine-tuning of both LLMs and PLP-former enhances the performance. Considering the computational cost, we adopt the fastest two-stage separate training strategy.
4.4 Limitation and Discussion
ProtChatGPT leverages the capabilities of LLMs for protein-specialized conversations. However, it inherits LLM’s potential language hallucination. It is an indispensable concern especially when it relates to protein research and healthcare. Although the two case studies in Section 4.3 demonstrate that our ProtChatGPT can distinguish some complex cases (e.g., homologous proteins and mutually exclusive functions), these capabilities heavily depend on the presence of similar examples in our training dataset, especially the description part. Given an unknown protein, ProtChatGPT might produce certain descriptions that sound correct but lack proper scientific verification, possibly leading researchers astray. This issue might be alleviated by training the model with more high-quality, aligned protein-text pairs, or aligning with more advanced LLMs in the future. In this manner, rigorous data processing and selection strategies should be implemented to ensure the validity and reliability of the training data. Feedback from domain experts is also important to refine the model. Combined with reinforcement or continual learning techniques, ProtChatGPT could keep improving the quality of its responses. Furthermore, training only one projection adapter might not provide enough capacity to learn extensive protein-text alignment. This issue could be alleviated by designing a more powerful adapter to facilitate the interactions between sequence and structure embeddings, or utilizing other powerful protein encoders such as GearNet (Zhang et al., 2023) and ESM-2 (Lin et al., 2023). In future work, addressing these issues and refining ProtChatGPT is essential. With ongoing improvements and regular expert feedback, ProtChatGPT has the potential to become a trusted assistant in protein research, offering more valuable insights for further investigations.
5 Conclusion
In this paper, we introduce ProtChatGPT, an AI-based protein chat system to implement ChatGPT-like functionalities for the protein research field. ProtChatGPT marks the initial effort at bootstrapping Protein-Language Pre-training (PLP) from pre-trained LLMs for various protein-language tasks. PLP sits at the intersection between protein and language, which effectively and efficiently enables ProtChatGPT to harvest the off-the-shelf large models from both protein and natural language communities. Experiments suggest that ProtChatGPT holds potential for application in protein understanding and design. We hope this work can facilitate protein research and further inspire other scientific disciplines.
Appendix A Details on Protein-Language Representation Learning
Obtaining queries that can extract informative protein representation regarding text is significant for protein-language alignment. To achieve this, we connect our PLP-former with the ESM-1b (Rives et al., 2021) model during the representation learning phase and train with the protein-language pairs. Following (Li et al., 2022a, 2023), we jointly train our model with three distinct pre-training tasks: Protein-Text Contrastive learning (PTC), Protein-grounded Text Generation (PTG), and Protein-Text Matching (PTM). Although these tasks utilize the same model structure and input format, they differ in the attention masking strategy applied between queries and text, thereby modulating their interaction.
Protein-Text Contrastive Learning (PTC).
For Protein-Text Contrastive Learning, by maximizing the mutual information, we aim to ensure the latent representation of protein and text are well-aligned. Specifically, given the query representation from the protein transformer $\boldsymbol{E}_{seq}$, we align it with the corresponding text embedding $t$. This is achieved by maximizing the similarity of positive pairs against those negative pairs where we directly use the embedding of $[cls]$ tokens from the text transformer as $t$. Given that the output of the protein transformer comprises multiple embeddings, we calculate the pairwise similarity between each query output and $t$. We then choose the highest value to represent the protein-text similarity. To prevent any information leakage, we utilize an unimodal self-attention mask that restricts direct interaction between queries and text.
Protein-grounded Text Generation (PTG).
The PTG task is designed to ensure that the learned queries can efficiently derive text-relevant information from the protein sequence. To accomplish this, we train the PLP-former to produce descriptions matching the respective protein sequences. Since the PLP-former prevents direct interaction between the frozen ESM-1b and text tokens, the data needed for description generation must first be garnered by the queries, ensuring efficient information extraction. We use a multimodal causal self-attention mask, to manage the interaction between queries and text. While queries can interact with one another, they cannot engage with the text tokens. Conversely, each text token can reference all queries as well as its preceding text tokens. Additionally, we substitute the $[CLS]$ token with a $[DEC]$ token at the beginning of the text sequence to indicate the decoding task.
Protein-Text Matching (PTM).
Protein-Text Matching task is leveraged for fine-grained protein-text representation alignment. This task is designed as a binary classification task where the model needs to determine if a given image-text pair aligns (positive) or misaligns (negative). We employ a bi-directional self-attention mask, allowing all queries and texts to mutually attend. As a result, the obtained query embeddings, $\boldsymbol{E}_{seq}$, encompass multimodal information. Each of these embeddings is then passed through a binary linear classifier to derive a logit, with the final matching score being the average of logits across all queries. For crafting informative negative pairs, we utilize the hard negative mining technique as described in (Li et al., 2021).
Appendix B Details and Examples of the Training Datasets
ProtDescribe Dataset. ProtDescribe dataset (Xu et al., 2023) is designed to augment protein sequences with text descriptions of their functions and other important properties. This dataset contains 549,000 pairs of proteins, including the EntryName, ProteinName, Function (obtained from Uniport), SubcellularLocation, Similarity, and ProteinSequence. In our implementation, we only use the protein sequence and the corresponding protein functions for the training of PLP-former. The PLP-former takes both sequence and descriptions as inputs, after three combined training tasks (Appendix. A, then outputs 32 learned tokens indicating the sequence embeddings that are most relevant to the descriptions. Some examples of this dataset are given in Table. 3.
Table 2: Examples of RCSB-PDB dataset with the protein entry id from Protein Data Bank.
EntryName ProteinName Function SubcellularLocation Similarity Sequence
14KL_BRUSI Lectin-like protein BA14k
Has immunoglobulin-binding and
hemagglutination properties,
and can bind to mannose.
Essential for virulence. May be
involved in LPS biosynthesis
or polysaccharide transport.
Cell membrane;
Single-pass
membrane protein
belongs to the BA14k family.
MNSFRKTCAGALA
LIFGATSIVPTVAAP
MNMDRPAINQNVI
QARAHYRPQNYNR
GHRPGYWHGHRG
YRHYRHGYRRHND
GWWYPLAAFGAGA
IIGGAISQPRPVYRAP
AG SPHVQWCYSRYK
SYRASDNTFQPYNGP
RKQCRSPYSR
11013_ASFWA Protein MGF 110-13L
Plays a role in virus cell tropism,
and may be required for efficient
virus replication in macrophages.
Host membrane;
Multi-pass
membrane protein
Belongs to the asfivirus MGF 110 family.
MGGGDYWPIIIRHCC
FYLVFSIAFVGYIVFA
YYKNLHLNTTMKLIA
LLCILIWLSQPGLNRP
LSIFYMKQNLPRTYTP
PIRELEYWCTYGKHC
DFCWECRNGICKNK
VWDDMPLIKQNDYIS
QCSIARYFDRCMYFIK
PKTPYIHYMDCSQPT
AYKGFSH
Table 3: Examples of ProtDescribe dataset.
Protein ID Abstract of the Primary Publication
1N21
The x-ray crystal structure of dimeric (+)-bornyl diphosphate synthase, a metal-requiring monoterpene cyclase from Salvia officinalis, is reported at
2.0-A resolution. Each monomer contains two alpha-helical domains: the C-terminal domain catalyzes the cyclization of geranyl diphosphate,
orienting and stabilizing multiple reactive carbocation intermediates; the N-terminal domain has no clearly defined function, although its N
terminus caps the active site in the C-terminal domain during catalysis. Structures of complexes with aza analogs of substrate and carbocation
intermediates, as well as complexes with pyrophosphate and bornyl diphosphate, provide "snapshots" of the terpene cyclization cascade.
2J11
The role of hydrophobic amino acids in the formation of hydrophobic cores as one of the major driving forces in protein folding has been extensively
studied. However, the implication of neutral solvent-exposed amino acids is less clear and available information is scarce. We have used a combinatorial
approach to study the structural relevance of three solvent-exposed residues (Tyr(327), Thr(329), and Gln(331)) located in thebeta-sheet of the
tetramerization domain of the tumor suppressor p53 (p53TD). A conformationally defined peptide library was designed where these three positions
were randomized. The library was screened for tetramer stability. A set of p53TD mutants containing putative stabilizing or destabilizing residue
combinations was synthesized for a thermodynamic characterization. Unfolding experiments showed a wide range of stabilities, with T(m) values
between 27 and 83 degrees C. Wild type p53TD and some highly destabilized and stabilized mutants were further characterized. Thermodynamic
and biophysical data indicated that these proteins were folded tetramers, with the same overall structure, in equilibrium with unfolded monomers.
An NMR study confirmed that the main structural features of p53TD are conserved in all the mutants analyzed. The thermodynamic stability of
the different p53TD mutants showed a strong correlation with parameters that favor formation and stabilization of the beta-sheet. We propose
that stabilization through hydrophobic interactions of key secondary structure elements might be the underlying mechanism for the strong influence
of solvent-exposed residues in the stability of p53TD.
3A08
Two crystal modifications of a collagen model peptide, (Pro-Pro-Gly)(4)-Hyp-Hyp-Gly-(Pro-Pro-Gly)(4) [where Hyp is (4R,2S)-L-hydroxyproline],
showed very similar unit-cell parameters and belonged to the same space group P2(1). Both crystals exhibited pseudo-merohedral twinning. The
main difference was in their molecular-packing arrangements. One modification showed pseudo-hexagonal packing, while the other showed
pseudo-tetragonal packing. Despite their different packing arrangements, no significant differences were observed in the hydration states of these
modifications. The peptide in the pseudo-tetragonal crystal showed a cyclic fluctuation of helical twists with a period of 20 A, while that in the
pseudo-hexagonal crystal did not. In these modifications, the puckering conformations of four of the 12 Hyp residues at the X position of the
Hyp(X)-Hyp(Y)-Gly sequence were in the opposite conformations to the previous hypothesis that Hyp(X) residues involved in Hyp(X):Hyp(Y)
and Hyp(X):Pro(Y) stacking pairs prefer up-puckering and down-puckering conformations, respectively. Detailed investigation of the molecular
interactions between Hyp(X) and adjacent molecules revealed that these opposite conformations appeared because the puckering conformation,
which follows the hypothesis, is subject to steric hindrance from the adjacent molecule.
RCSB-PDB Dataset. RCSB-PDB dataset (Guo et al., 2023) is originally sourced from the Research Collaboratory for Structural Bioinformatics Protein Data Bank  [^1] https://www.rcsb.org, which includes 204,826 experimentally determined 3D structures. ProteinChat (Guo et al., 2023) further selected 143,508 proteins that have a primary publication linked with a PubMed ID and can be extracted with a valid chain. In our implementation, we compile both 1D amino acid sequences and 3D atomic coordinates, along with the abstract of corresponding scientific literature for the second-stage training. With the PLP-former and structure encoder frozen, LLM takes out the sequence embedding, structure embedding and the user question as inputs, and produces the related answers. Some examples of this dataset are given in Table. 2.
Appendix C Details on Metrics
In our implementation, we use seven different metrics on the proposed ProtChatGPT to verify the performance. These metrics not only focus on the low-level lexical and syntactic alignment but also on high-level semantic information.
BLEU (Papineni et al., 2002) (Bilingual Evaluation Understudy) serves as a crucial metric for assessing the quality of the machine-generated text, particularly in machine translation contexts. It quantifies the similarity between the candidate and reference text, yielding a score within the range of 0 to 1. A higher BLEU score indicates a closer match between the candidate and reference texts. BLEU is mathematically defined as follows:
$p_{n}=\frac{\sum\limits_{C\in\{Candidates\}}\sum\limits_{\textit{n-gram}\in C}Count_{clip}(\textit{n-gram})}{\sum\limits_{C^{\prime}\in\{Candidates\}}\sum\limits_{\textit{n-gram}^{\prime}\in C^{\prime}}Count(\textit{n-gram}^{\prime})}$ (6)
$\text{BP}=\begin{cases}1&\text{if }c>r\\ e^{(1-r/c)}&\text{if }c\leq r\end{cases}$ (7)
$\text{BLEU}=\text{BP}\cdot exp\left(\sum_{n=1}^{N}w_{n}logp_{n}\right)$ (8)
where BP (Brevity Penalty) helps penalize overly short translations and $p_{n}$ represents the precision of n-grams, ranging from 1 to a predefined maximum.
ROUGE-L (Lin & Hovy, 2002) finds frequent use in the automatic evaluation of text summarization and machine translation. It calculates the longest common subsequence between the candidate and reference texts, prioritizing recall over precision. ROUGE-L can be expressed mathematically as:
$R_{lcs}=\frac{LCS(X,Y)}{m}$ (9)
$P_{lcs}=\frac{LCS(X,Y)}{n}$ (10)
[…]
In this context, $X$ represents the predicted text with a length of $n$. $Y$ represents the ground truth text with a length of $m$. $\beta$ is a hyperparameter used to adjust the emphasis on precision and recall. LCS calculates the length of the longest common subsequence, $R_{lcs}$ measures recall, and $P_{lcs}$ measures precision, respectively.
METEOR (Banerjee & Lavie, 2005) provides a comprehensive evaluation of machine-generated text by considering not only exact word matches but also synonyms and stemming. It combines precision, recall, and alignment factors to offer a holistic assessment. METEOR is mathematically represented as:
$F=\frac{(\alpha^{2}+1)P}{R+\alpha P}$ (12)
$\text{Meteor}=(1-Penalty)\cdot F$ (13)
where $Penalty$ penalizes excessive word mismatches, $\alpha$ is a configurable parameter, $R$ and $P$ represents recall and precision respectively.
CIDEr (Vedantam et al., 2015) (Consensus-based Image Description Evaluation) primarily assesses the quality of image captions produced by automated systems. It places importance on consensus among multiple reference captions and emphasizes the inclusion of diverse descriptive words.
The mathematical formulation of CIDEr is given by:
$CIDEr_{n}(c,S)=\frac{1}{M}\sum_{i=1}^{M}\frac{g^{n}(c)\cdot g^{n}(S_{i})}{||g^{n}(c)||\times||g^{n}(S_{i})||}$ (14)
In the formula, $c$ represents the candidate text, $S$ denotes the set of reference texts, $n$ specifies the use of n-grams, $M$ represents the number of reference texts, and $g$ corresponds to the TF-IDF vector based on n-grams.
SPICE (Anderson et al., 2016) offers a metric designed to evaluate the semantic content of image captions, with a focus on their precision within generated captions.
SPICE’s mathematical expression is as follows:
$P(c,S)=\frac{|T(G(c))\otimes T(G(S))|}{|T(G(c))|}$ (15)
$R(c,S)=\frac{|T(G(c))\otimes T(G(S))|}{|T(G(S))|}$ (16)
$SPICE(c,S)=\frac{2\cdot P(c,S)\cdot R(c,S)}{P(c,S)+R(c,S)}$ (17)
where the binary matching operator $\otimes$ is the function that returns matching tuples in two scene graphs, $P$ represents the precision of semantic propositions and $R$ signifies the recall of semantic propositions, respectively.
BertScore (Zhang et al., 2019) is a metric that leverages contextual embeddings from BERT models to assess the quality of machine-generated text. It measures the similarity between the candidate text and the reference text using contextual embeddings.
PubMed BERTScore comes from the classical BertScore (Zhang et al., 2019), which is a metric that leverages contextual embeddings from BERT models to assess the quality of machine-generated text. It measures the similarity between the candidate text and the reference text using contextual embeddings. In our implementation, to better assess the quality of ProtChatGPT in generating protein-related descriptions, we further replaced the encoder with the encoder of PubMedBERT (Gu et al., 2021). PubMedBERT is the latest BERT (Devlin et al., 2018) model pre-trained on the biomedical corpus, which outperformed BioBERT on the BLURB (Gu et al., 2021) (Biomedical Language Understanding and Reasoning Benchmark).